# Very High-order Compact Gas-kinetic Scheme With Discontinuity Feedback Factor

### v1

@@ Abstract
This paper presents a robust and efficient very high-order scheme for compressible flow simulation, addressing critical limitations of existing high-order methods. The proposed scheme combines the compact gas-kinetic scheme (CGKS) with an adaptive stencil extension reconstruction with discontinuity feedback factor (ASE-DFF), achieving significant improvements in both robustness and computational efficiency. Traditional weighted essentially non-oscillatory (WENO) schemes suffer from reduced robustness at higher order and require costly smoothness indicators for large stencils. Meanwhile, compact methods based on Discontinuous Galerkin (DG) and Flux Reconstruction (FR) struggle with poor time-marching efficiency. In contrast, the ASE-DFF-CGKS introduces two key innovations: (1) a unified framework enabling arbitrarily high-order compact gas-kinetic scheme without sacrificing large CFL number, and (2) a discontinuity feedback factor that eliminates the need for expensive smoothness indicator calculations while essentially keeping first-order robustness near discontinuities. The scheme’s advantages are demonstrated through benchmark simulations. It maintains a CFL number above 0.5 for up to 9th-order case, unlike conventional compact methods that restrict a CFL less than 0.05. Also it delivers high-resolution results for flow involving strong shock and rarefaction wave. This work provides a practically impactful solution for high-fidelity compressible flow simulation, balancing computational efficiency, high-order accuracy and robustness in challenging flow regimes.
@@
compact gas-kinetic scheme, very high-order, discontinuity feedback factor, adaptive stencil extension reconstruction
@@ 1 Introduction
Over the past few decades, significant advancements have been made in high-order computational fluid dynamics (CFD) methods for solving the Euler and Navier-Stokes equations in compressible flow. Among these, weighted essentially non-oscillatory (WENO) scheme [16, 24] on structured meshes have gained prominence due to their simplicity, robustness and ability to achieve high-order accuracy [1]. However when applied to unstructured mesh, WENO scheme face challenges related to large stencil, which complicate their implementation in complex geometries. In contrast, high-order compact methods such as discontinuous Galerkin (DG) [5, 6] flux reconstruction (FR) [32] and correction procedure via reconstruction (CPR) [26, 33] frameworks offer superior geometric flexibility and efficient parallelization. Despite these advantages, their stringent CFL condition impose severe restrictions on the allowable time step, limiting computational efficiency. Furthermore, these methods still encounter difficulties in accurately capturing shock discontinuities. To address these limitations, the compact gas-kinetic scheme (CGKS) [11, 43, 14, 15, 12] has emerged as an effective alternative. By combining a two-stage fourth-order gas-kinetic scheme with a compact Hermite WENO reconstruction, CGKS achieves high accuracy and resolution while maintaining robust performance. Notably, CGKS permits larger CFL number (typically around 0.5 or higher), enhancing computational efficiency without sacrificing stability. Its success in simulating compressible flows with strong shock discontinuity underscores its potential as a reliable tool for complex flow simulation.
@@
The compact gas-kinetic scheme (CGKS) [11, 43, 22] represents an innovative approach in computational fluid dynamics. This method demonstrates remarkable capability in achieving third-order accuracy within a compact framework on three-dimensional unstructured mesh [14, 15, 12]. Comprehensive numerical experiments involving complex three-dimensional compressible flow have consistently validated both its robustness and accuracy [38, 39]. A fundamental distinction exists between CGKS and conventional CFD methods. The method builds upon the classical two-stage fourth-order GKS formulation introduced by Pan and Xu [27]. Departing from the widely-used fourth-order Runge-Kutta (RK4) temporal discretization [10], the scheme incorporates concepts from the two-stage fourth-order GRP approach developed by Li [21]. This implementation capitalizes on the time-dependent gas-kinetic flux distribution function. Spatial discretization employs traditional WENO to determine conservative variable for the gas-kinetic flux function at cell interface. Recognizing the inherent complexity of gas-kinetic flux evaluation, Ji et al. have proposed enhanced variants including WENO-AO GKS to streamline computations [37, 13]. While standard high-order gas-kinetic scheme with WENO reconstruction underutilize the method’s full potential, CGKS effectively combines three key elements including GKS flux formulation, two-stage fourth-order temporal discretization and Hermite WENO reconstruction. The solution evolution process in CGKS shares conceptual similarities with discontinuous Galerkin (DG) methods and HWENO schemes [28], requiring simultaneous updates of both conservative variables and their gradients within each computational cell. The relative simplicity of the CGKS framework stems principally from unique properties of the gas-kinetic flux. By fully exploiting the spatiotemporal evolution characteristics of the gas-kinetic distribution function at interfaces, CGKS enables instantaneous determination of conservative variables through moment integration of the distribution function. This elegant formulation permits straightforward gradient updates via Gauss’s theorem with notable efficiency.
@@
The WENO scheme has undergone extensive development and found widespread application in computational fluid dynamics. In the context of high-speed turbulence simulations, the direct numerical simulation (DNS)  [20, 8]approach predominantly employs high-order finite difference method incorporating WENO scheme. Subsequent advancements by Qiu and Zhu introduced the Hermite WENO scheme [28, 23, 44], which utilizes gradient information of conservative variables within each cell. Similar to conventional WENO, this method constructs high-order linear polynomials using cell-centered information and employs smoothness indicators to weight sub-stencils for discontinuity resolution. The key distinction lies in its ability to achieve enhanced accuracy on more compact stencils through the incorporation of gradient data. From a physical perspective, these compact stencils demonstrate superior performance in resolving discontinuities across computational domains. To better integrate WENO within finite volume framework, the development of WENO-AO scheme has successfully addressed issues related to negative weights during Gaussian point value reconstruction [1, 13]. Recent progress in the discontinuity feedback factor [12], which has shown remarkable success when implemented in high-order numerical method [12, 40]. This innovation has exhibited exceptional robustness and shock-capturing capability in three-dimensional supersonic and hypersonic flow simulation. Further refinements led to the development of adaptive stencil extension with discontinuous feedback reconstruction on structured mesh [41], maintaining solution robustness while efficiently achieving seventh-order and ninth-order accuracy through non-compact stencil combined with WENO scheme. This advancement proves particularly valuable as the computation of smoothness indicators for high-order stencils typically incurs significant computational expense. Consequently, the present work employs the discontinuity feedback factor to realize higher-order spatial accuracy within the compact gas-kinetic scheme framework [11, 43].The compact gas-kinetic scheme with discontinuity feedback factor proposed in this paper, achieves 7th-order and 9th-order spatial accuracy using a central stencil. Compared with the non-compact high-order gas-kinetic scheme, it uses the same stencil size as the 5th-order non-compact HGKS. And it overcomes the main drawbacks of the original compact GKS, including the 6th-order and 8th-order compact GKS. Early the CGKS constructed the non-equilibrium and equilibrium state on both sides of the interface separately. But subsequent non-compact and compact HGKS have abandoned this approach, directly using kinetic weighting of the non-equilibrium state to obtain the equilibrium state. The disadvantage of larger stencil selection is that constructing the equilibrium state on both sides of the interface still requires a non-compact stencil. By employing the discontinuity feedback factor strategy, a compact stencil can be used to construct the 5th-order, 7th-order and 9th-order CGKS scheme, obtaining the equilibrium state by kinetic weighting of higher-order non-equilibrium state. The tangential reconstruction in the two-dimensional reconstruction process of the CGKS based on the discontinuity feedback factor is also simplified. Moreover, the high-order CGKS based on the discontinuity feedback factor can avoid the computation of high-order smoothness indicators, thereby reducing the computational cost. The discontinuity feedback factor plays a crucial role in shock capturing. Accurate shock discontinuity capture and compression of the slopes of higher-order reconstruction polynomials provide the CGKS with good robustness.
@@
The organization of this paper is as follows. The Section 2 mainly introduces the gas-kinetic scheme and the two-stage fourth-order time advancement method. The Section 3 mainly presents the discontinuity feedback factor and a series of compact stencils for achieving high-order accuracy in the compact gas-kinetic scheme. Additionally, the two-dimensional reconstruction framework of the CGKS is given in detail. The Section 4 proves the performance such as the accuracy and robustness of the scheme through numerical cases. In conclusion, a summary of the current work is provided.
@@ 3.3 Two dimensional reconstruction
Before introducing the reconstruction procedure, let’s denote $\overline{W}$ as cell averaged, $\hat{W}$ as line averaged, and $W$ as pointwise values. Here $W^{l,r}$ represent the reconstructed quantities on the left and right sides, which correspond to the non-equilibrium initial part in GKS framework.
At $t^{n}$ step, for cell $(i,j)$ the cell average quantities $\overline{W}_{i,j}^{n}$; $\overline{W}_{x,i,j}^{n}$, $\overline{W}_{y,i,j}^{n}$ are stored. For a fifth-order scheme, two Gaussian points are required at each interface for numerical flux integration. For the seventh-order scheme, three Gaussian points are needed, and for the ninth-order scheme, four Gaussian points are required. Our target is to construct $W^{l,r}_{x},W^{l,r}_{x},W^{l,r}_{xx},W^{l,r}_{y},W^{l,r}_{yy},W^{l,r}_{xy}$ at each Gaussian point. To obtain these quantities, the line averaged slopes $\hat{W}^{n}_{x,i,j_{l}},\hat{W}^{n}_{y,i_{l},j}$ are additionally evaluated, where l=1,,ml=1,...,m corresponds to the number of Gaussian points required at each interface. For a better illustration, a schematic is plotted in Fig. 2 and the reconstruction procedure for the Gaussian point $i+1/2,j_{l}$ is summarized as follows. Here the time level $n$ is omitted.
@@ 4.5 Low density and low pressure problems
In Fig. 8-9, the density, velocity and pressure distribution of two numerical tests are shown for ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS. The results of the three schemes are in good agreement with the reference solutions. In the results of the double rarefaction wave problem and the Le Blanc problem, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS have higher resolution compared to the 5-th order ASE-DFF(5,3)-CGKS. Overall, the discontinuity feedback factor is good at simulating some tests involving strong shock waves and strong rarefaction waves. Therefore, the compact GKS using the discontinuity feedback factor can maintain good robustness in extreme case.
@@ 4.7 Two-dimensional Riemann problems
The mesh is $500\times 500.$ The results of the Configuration 3 at $t=0.6$ for the ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS, and ASE-DFF(9,7,5,3)-CGKS are presented in Fig. 11.The proof shows that the adaptive high-order CGKS based on discontinuity feedback factor demonstrates extremely high resolution. The three schemes have revealed numerous small-scale structures. As the order increases, the performance continuously improves.
@@ 4.8 Double Mach reflection problem
The density distributions are shown in Fig. 12. In the double Mach reflection problem, it is proved that all three compact CGKS schemes have excellent robustness and shock-capturing capability. At the same time, ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS all have lower dissipation. They can capture a larger number of small scale flow structures. However, as higher-order compact linear stencils are added to the reconstruction, some numerical oscillations can also be observed to increase.
Figure 12: The density distributions for Double Mach reflection problem: ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS scheme. Mesh: $960\times 240.$ $CFL=0.5.$ $T=0.2.$

### v2

@@ Abstract
This paper presents a robust and efficient very high-order scheme for compressible flow simulation, addressing key limitations of existing high-order methods. The proposed approach integrates the compact gas-kinetic scheme (CGKS) with an adaptive stencil extension reconstruction with a discontinuity feedback factor (ASE-DFF), resulting in significant improvements in both robustness and computational efficiency. Traditional weighted essentially non-oscillatory (WENO) schemes often exhibit reduced robustness at higher orders and require computationally expensive smoothness indicators for large stencils. On the other hand, compact methods based on Discontinuous Galerkin (DG) and Flux Reconstruction (FR) suffer from low time-marching efficiency. In contrast, the ASE-DFF-CGKS introduces two major innovations: (1) a unified framework that enables arbitrarily high-order compact gas-kinetic schemes without sacrificing a large CFL number, and (2) a discontinuity feedback factor that eliminates the need for costly smoothness indicator calculations while maintaining first-order robustness near discontinuities. The advantages of the proposed scheme are demonstrated through benchmark simulations, where it sustains a CFL number above 0.5 for schemes up to ninth order—significantly higher than conventional compact methods. Moreover, it delivers high-resolution results for flows involving strong shocks and rarefaction waves. Overall, this work offers a practically impactful solution for high-fidelity compressible flow simulation, achieving an effective balance between computational efficiency, high-order accuracy, and robustness in challenging flow regimes.
@@
compact gas-kinetic scheme, very high-order methods, discontinuity feedback factor, adaptive stencil extension reconstruction
@@ 1 Introduction
Over the past few decades, high-order computational fluid dynamics (CFD) methods for solving the Euler and Navier-Stokes equations in compressible flows have seen significant advancements. Among these methods, weighted essentially non-oscillatory (WENO) schemes [20, 29] on structured meshes have gained prominence for their simplicity, robustness, and ability to achieve high-order accuracy [3]. However, when applied to unstructured meshes, WENO schemes face challenges associated with large stencils, which complicate their implementation in complex geometries. In contrast, high-order compact methods such as discontinuous Galerkin (DG)[8, 9], flux reconstruction (FR)[39], and correction procedure via reconstruction (CPR) [31, 40] frameworks offer greater geometric flexibility and efficient parallelization. Despite these advantages, their stringent CFL conditions impose severe restrictions on the allowable time step, thus limiting computational efficiency. Additionally, these methods continue to face challenges in accurately capturing shock discontinuities. To address these limitations, the compact gas-kinetic scheme (CGKS) [15, 50, 18, 19, 16] has been developed, enabling the use of larger CFL numbers (typically around 0.5 or higher) while stably resolving strong shocks in complex compressible flows.
@@
The compact gas-kinetic scheme (CGKS)[15, 50, 27] represents an innovative advancement in computational fluid dynamics. This method demonstrates exceptional capability in achieving third-order accuracy within a compact framework on three-dimensional unstructured meshes[18, 19, 16]. Extensive numerical experiments involving complex three-dimensional compressible flows have consistently validated its robustness and accuracy [45, 46]. A fundamental distinction sets CGKS apart from conventional CFD methods. Building upon the classical two-stage fourth-order gas-kinetic scheme (GKS) formulation introduced by Pan and Xu [32], CGKS departs from the widely used fourth-order Runge-Kutta (RK4) temporal discretization [14], instead incorporating concepts from the two-stage fourth-order generalized Riemann problem (GRP) approach developed by Li [26]. This implementation leverages the time-dependent gas-kinetic flux distribution function, while spatial discretization employs traditional WENO reconstruction to determine conservative variables at cell interfaces for the gas-kinetic flux function. Recognizing the inherent complexity of gas-kinetic flux evaluation, Ji et al. have proposed enhanced variants, such as the WENO-AO GKS, to streamline computations [44, 17]. While standard high-order gas-kinetic schemes with WENO reconstruction may underutilize the method’s full potential, CGKS effectively combines three key elements: the GKS flux formulation, two-stage fourth-order temporal discretization, and Hermite WENO reconstruction. The solution evolution process in CGKS shares conceptual similarities with discontinuous Galerkin (DG) methods and Hermite WENO (HWENO) schemes [35], requiring simultaneous updates of both conservative variables and their gradients within each computational cell. The relative simplicity of the CGKS framework arises primarily from the unique properties of the gas-kinetic flux. By fully exploiting the spatiotemporal evolution characteristics of the gas-kinetic distribution function at interfaces, CGKS enables the instantaneous determination of conservative variables through moment integration of the distribution function. This elegant formulation also permits straightforward gradient updates via Gauss’s theorem, offering notable computational efficiency.
@@
The WENO scheme has undergone extensive development and has found widespread application in computational fluid dynamics. For high-speed turbulence simulations, direct numerical simulation (DNS) approaches [25, 12] predominantly employ high-order finite difference methods that incorporate WENO schemes. Building on this foundation, Qiu and Zhu introduced the Hermite WENO scheme [35, 28, 51], which utilizes the gradient information of conservative variables within each cell. Similar to conventional WENO methods, Hermite WENO constructs high-order polynomials using cell-centered information and employs smoothness indicators to weight sub-stencils for effective discontinuity resolution. The key distinction lies in its ability to achieve enhanced accuracy on more compact stencils by incorporating gradient data. Physically, these compact stencils demonstrate superior performance in resolving discontinuities across computational domains. To facilitate the integration of WENO schemes within the finite volume framework, the development of the WENO-AO scheme has effectively addressed issues related to negative weights during Gaussian point value reconstruction [3, 17]. Recent advances involving the discontinuity feedback factor [16] have demonstrated remarkable success when implemented in high-order numerical methods [16, 47], exhibiting exceptional robustness and shock-capturing capability in three-dimensional supersonic and hypersonic flow simulations. Further refinements have led to the development of adaptive stencil extension with discontinuity feedback reconstruction on structured meshes [48], which maintains solution robustness while efficiently achieving seventh- and ninth-order accuracy through non-compact stencils combined with the WENO scheme. This advancement is particularly valuable, as the computation of smoothness indicators for high-order stencils typically entails significant computational expense.
For high-order schemes, it is crucial to design methods capable of maintaining high accuracy in the presence of shock waves or turbulence [34]. Achieving seventh-, ninth-, or even higher-order accuracy is essential for resolving flows that contain shocks or turbulence. However, increasing the order of accuracy inevitably poses severe challenges to robustness. The primary reason is that higher-order reconstructions substantially reduce numerical dissipation, making it more difficult to suppress non-physical oscillations near shocks or strong discontinuities [4, 38]. The typical solution involves optimizing the nonlinear stability mechanism and appropriately reducing the order at discontinuity points. For WENO and TENO schemes, adaptive-accuracy variants such as WENO-AO [3], TENO-AA [11], and other hybrid WENO schemes [1, 33] can deliver ninth- or tenth-order accuracy. However, beyond the significant computational cost of evaluating smoothness indicators [2], these schemes often require complex stencil-based smoothness assessment mechanisms or rely on positivity-preserving fluxes to pass stringent benchmark cases. In the case of DG formulations, the incorporation of limiters or artificial viscosity is necessary, yet designing high-order limiters remains notoriously challenging and often leads to excessive loss of accuracy [22]. Moreover, the large stencils required by very high-order schemes complicate the imposition of boundary conditions and limit adaptability to unstructured grids, thereby undermining robustness.
This work employs the discontinuity feedback factor to achieve higher-order spatial accuracy within the compact gas-kinetic scheme (CGKS) framework [15, 50]. The proposed CGKS with a discontinuity feedback factor attains seventh- and ninth-order spatial accuracy using a central stencil. Compared to non-compact high-order gas-kinetic schemes, it utilizes the same stencil size as the fifth-order non-compact HGKS, while overcoming the primary limitations of previous compact GKS methods, including sixth- and eighth-order compact GKS. Earlier CGKS formulations constructed the non-equilibrium and equilibrium states separately on both sides of the interface. However, both subsequent non-compact and compact HGKS abandoned this approach, instead obtaining the equilibrium state directly through kinetic weighting of the non-equilibrium state. The drawback of using larger stencils is that equilibrium state construction at the interface still requires a non-compact stencil. By introducing the discontinuity feedback factor strategy, it becomes possible to construct fifth-, seventh-, and ninth-order CGKS using compact stencils, obtaining the equilibrium state through kinetic weighting of higher-order non-equilibrium states. In addition, the tangential reconstruction process in the two-dimensional CGKS is simplified when employing the discontinuity feedback factor. Furthermore, the high-order CGKS based on this strategy eliminates the need for computationally expensive high-order smoothness indicators, thereby reducing overall computational cost. The discontinuity feedback factor plays a crucial role in shock capturing, enabling accurate resolution of shock discontinuities and effective compression of higher-order reconstruction polynomial slopes, thus imparting strong robustness to the CGKS.
@@
The organization of this paper is as follows: Section 2 introduces the gas-kinetic scheme and the two-stage fourth-order time advancement method. Section 3 presents the discontinuity feedback factor and a series of compact stencils for achieving high-order accuracy in the compact gas-kinetic scheme; the two-dimensional reconstruction framework of the CGKS is also detailed. Section 4 demonstrates the performance of the scheme, including its accuracy and robustness, through various numerical cases. Finally, the paper concludes with a summary of the current work.
@@ 3.2 Adaptive stencil extension reconstruction
Remark 1. We assess the computational complexity of r-th-order reconstruction for the conventional WENO reconstruction and the ASE-DFF reconstruction. In WENO/HWENO formulations the reconstruction polynomial is assembled as the nonlinear convex combination of the pointwise values obtained from $\mathcal{O}(r)$ distinct sub-stencils. Each sub-stencil requires $\mathcal{O}(r)$ operations, and exactly $\mathcal{O}(r)$ such polynomials are combined. So the computational complexity of the polynomial calculation is $\mathcal{O}(r^{2})$. Smoothness indicators are evaluated on the same sub-stencils. Each sub-stencil corresponds to a polynomial of $(r+1)/2$-th order . Hence, according to Eq.(43), up to (r1)/2(r-1)/2 successive derivatives must be computed. This leads to an overall computational complexity of $\mathcal{O}(r^{3})$ for the smoothness indicators evaluation. Consequently, the total computational complexity of the classical WENO/HWENO reconstruction is $\mathcal{O}(r^{3})$. By contrast, the ASE-DFF reconstruction attains the r-th order by directly finishing the high-order reconstruction with the corresponding DFF. The DFF itself can be inherited and no smoothness indicators are required. The computational complexity of the ASE-DFF reconstruction is only $\mathcal{O}(r)$. Therefore, the cost of both computation and memory for ASE-DFF method are relatively small.
@@ 3.3 Two dimensional reconstruction
Before introducing the reconstruction procedure, let’s denote $\overline{W}$ as cell averaged, $\hat{W}$ as line averaged, and $W$ as pointwise values. Here $W^{l,r}$ represent the reconstructed quantities on the left and right sides, which correspond to the non-equilibrium initial part in GKS framework. At $t^{n}$ step, for cell $(i,j)$ the cell average quantities $\overline{W}_{i,j}^{n}$; $\overline{W}_{x,i,j}^{n}$, $\overline{W}_{y,i,j}^{n}$ are stored. For a fifth-order scheme, two Gaussian points are required at each interface for numerical flux integration. For the seventh-order scheme, three Gaussian points are needed, and for the ninth-order scheme, four Gaussian points are required. Our target is to construct $W^{l,r}_{x},W^{l,r}_{x},W^{l,r}_{xx},W^{l,r}_{y},W^{l,r}_{yy},W^{l,r}_{xy}$ at each Gaussian point. To obtain these quantities, the line averaged slopes $\hat{W}^{n}_{x,i,j_{l}},\hat{W}^{n}_{y,i_{l},j}$ are additionally evaluated, where l=1,,ml=1,...,m corresponds to the number of Gaussian points required at each interface. For a better illustration, a schematic is plotted in Fig. 2 and the reconstruction procedure for the Gaussian point $i+1/2,j_{l}$ is summarized as follows. Here the time level $n$ is omitted.
@@ 4.5 Low density and low pressure problems
In Fig. 8-9, the density, velocity and pressure distribution of two numerical tests are shown for ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS. The results of the three schemes are in good agreement with the reference solutions. In the results of the double rarefaction wave problem and the Le Blanc problem, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS have higher resolution compared to the 5-th order ASE-DFF(5,3)-CGKS. Overall, the discontinuity feedback factor is good at simulating some tests involving strong shock waves and strong rarefaction waves. Furthermore, only ASE-DFF-CGKS was able to pass the Le Blanc problem. The traditional WENO-GKS or HWENO-CGKS gas kinetic schemes fail to pass the Le Blanc problem test.Therefore, the compact GKS using the discontinuity feedback factor can maintain good robustness in extreme case.
@@ 4.7 Two-dimensional Riemann problems
The mesh is $500\times 500.$ The results of the Configuration 3 at $t=0.6$ for the ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS, and ASE-DFF(9,7,5,3)-CGKS are presented in Fig. 11.The proof shows that the adaptive high-order CGKS based on discontinuity feedback factor demonstrates extremely high resolution. Furthermore, the ASE-DFF-CGKS has better resolution than the HWENO-CGKS. The ASE-DFF-CGKS have revealed numerous small-scale structures. As the order increases, the performance continuously improves.
@@ 4.8 Double Mach reflection problem
The density distributions are shown in Fig. 12. In the double Mach reflection problem, it is proved that all three compact CGKS schemes have excellent robustness and shock-capturing capability. At the same time, ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS all have lower dissipation and better resolution than the HWENO-CGKS. They can capture a larger number of small scale flow structures. However, as higher-order compact linear stencils are added to the reconstruction, some numerical oscillations can also be observed to increase.
Figure 12: The density distributions for Double Mach reflection problem. Mesh: $960\times 240.$ $CFL=0.5.$ $T=0.2.$